2. Famous Cinema Theatres sell the movie tickets for $16 per ticket for evening screenings and sell 250 tickets a night on average. They estimate that the price elasticity of demand for tickets is (negative) 0.8.

Calculate the expected number of tickets sold if they reduce the ticket price to $14.
1
Expert's answer
2019-05-27T11:58:59-0400

Using the formula for price elasticity of demand:

Ed = (Q2 - 250)/(14 - 16)×(14 + 16)/(Q2 + 250) = -0.8,

(Q2 - 250)/2×30/(Q2 + 250) = 0.8,

0.8×(Q2 + 250) = 30×(Q2 - 250),

29.2×Q2 = 7700,

Q2 = 7700/29.2 = 263.7 units.